# DeepPresenter: Environment-Grounded Reflection for Agentic Presentation Generation

### v1

@@ 5.3 Efficiency Analysis
Figure 6 presents the cost-performance trade-off across frameworks and models. (1) DeepPresenter-9B advances the Pareto frontier, significantly outperforming the prior frontier point at comparable cost. Compared to KCTV + Gemini-3-Pro (3.83), DeepPresenter-9B achieves 4.19 at a similar price, a significant improvement in cost-quality trade-off. (2) DeepPresenter establishes a new upper bound for presentation generation, surpassing the previous best system Gamma. With an average score of 4.44 versus Gamma’s 4.36, DeepPresenter delivers the strongest result in our evaluation.
@@ 6 Related Work
Compared with previous methods, DeepPresenter formulates presentation generation as an autonomous exploration and collaboration process between two specialized agents. The Researcher-Presenter decomposition enables adaptive planning based on task complexity, while environment-grounded reflection allows agents to verify and refine artifacts through rendered slides and file system states (Tang et al., 2025; Jiang et al., 2025; Stechly et al., 2024).
@@ A.3 Tool Usage Analysis
Across domains (Figure 8b), Researcher shows adaptable usage patterns reflecting task characteristics. PersonaHub tasks exhibit significantly higher Retrieve usage, as persona-driven queries do not provide reference documents, requiring agents to actively search for relevant materials. In contrast, arXiv and FinePDF tasks involve provided source documents, leading to higher File usage for document processing and lower reliance on retrieval. Tool categories are detailed in Table 8.

### v2

@@ 5.3 Efficiency Analysis
Figure 6 presents the cost-performance [^2] The cost of fine-tuned models is estimated using the per-token pricing of their base models on OpenRouter. trade-off across frameworks and models. (1) DeepPresenter-9B advances the Pareto frontier, significantly outperforming the prior frontier point at comparable cost. Compared to KCTV + Gemini-3-Pro (3.83), DeepPresenter-9B achieves 4.19 at a similar price, a significant improvement in cost-efficiency. (2) DeepPresenter establishes a new upper bound for presentation generation, surpassing the previous best system Gamma. With an average score of 4.44 versus Gamma’s 4.36, DeepPresenter delivers the strongest result in our evaluation.
@@ 6 Related Work
Compared with previous methods, DeepPresenter formulates presentation generation as an autonomous exploration and collaboration process between two specialized agents. The Researcher-Presenter decomposition enables adaptive planning based on task complexity, while environment-grounded reflection allows agents to verify and refine artifacts through rendered slides and file system states (Tang et al., 2025; Jiang et al., 2025).
@@ 7 Conclusion
Acknowledgments
We sincerely thank the reviewers for their insightful comments and valuable suggestions. This work was supported by Beijing Natural Science Foundation (L243006), the Natural Science Foundation of China (No. 62476265, 62306303, 62506354).
@@ A.3 Tool Usage Analysis
Across domains (Figure 8b), Researcher shows adaptable usage patterns reflecting task characteristics. PersonaHub tasks exhibit significantly higher Retrieve usage, as persona-driven queries do not provide reference documents, requiring agents to search for relevant materials Yuan et al. (2025). In contrast, arXiv and FinePDF tasks involve provided source documents, leading to higher File usage for document processing and lower reliance on retrieval. Tool categories are detailed in Table 8.